OpenStudy (anonymous):

Evaluate cos(sin^-1(2x))

OpenStudy (anonymous):

evaluate??

OpenStudy (anonymous):



OpenStudy (anonymous):

by pythagoras you get \[\sqrt{1-(2x)^2}\] or \[\sqrt{1-4x^2}\] so \[\cos(\sin^{-1}(2x))=\sqrt{1-4x^2}\]

OpenStudy (anonymous):

I don't understand the way it was worked out.

OpenStudy (aum):

Find cos( sin^-1(2x) ) Assume sin^-1(2x) = t. We need to find cos(t). Since sin^-1(2x) = t, it implies sin(t) = 2x Now refer to the diagram that satellite drew above where the bottom left angle is t. The dimensions are so chosen in the triangle to make sin(t) = 2x sin(t) = opposite / hypotenuse = 2x / 1 = 2x Now we can easily find cosine of the angle t which is adjacent / hypotenuse. Find the adjacent using Pythagoras Theorem. Hypotenuse = 1. Plug the values and find cos(t).